OpenStudy (anonymous):

A sphere has a radius of 3x+2. Which polynomial in standard form best describes the total surface area of the sphere?

OpenStudy (anonymous):

Note: Use the formula S = 4r2 for the surface area of a sphere. don't pay attention to this if you don't want, I'm just thinking

OpenStudy (anonymous):

@crazyinlove any given answers?

OpenStudy (anonymous):

16\[\pi\] x^2+16\[\pi\] 16 pi x^2+48pi x+16 pi 36pix^2+48 pi x+16 pi 36 pi x^2+24 pi x+ 16 pi

OpenStudy (anonymous):

note: 4(3x+2)2

OpenStudy (anonymous):

ok im not completely sure but i think its "16 pi x^2+48pi x+16 pi"

OpenStudy (anonymous):

can u put the work up please

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

radius = 3x+2 diameter = (3x+2)*2 area = (Pi(3.14)*radius^2) = (3.14*(3x+2)^2) | PI(3x+2)2

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

np and thats as far as i can go... my mind can't handle it XD

OpenStudy (anonymous):

lol